I'm having trouble with my honors packet for high school. I'm really stuck with this question. X^2+2x-15=0. Thanks

Answer by algebrahouse.com(1659)

x² + 2x - 15 = 0
(x + 5)(x - 3) = 0 {factored into two binomials}
x + 5 = 0 or x - 3 = 0 {set each factor equal to 0}
x = -5 or x = 3 {solved each equation for x}
www.algebrahouse.com